# MTVQA: Benchmarking Multilingual Text-Centric Visual Question Answering

### v3

@@ Abstract
Text-Centric Visual Question Answering (TEC-VQA) in its proper format not only facilitates human-machine interaction in text-centric visual environments but also serves as a de facto gold proxy to evaluate AI models in the domain of text-centric scene understanding. Nonetheless, most existing TEC-VQA benchmarks focus on high-resource languages like English and Chinese. Despite pioneering works expanding multilingual QA pairs in non-text-centric VQA datasets through translation engines, the translation-based protocol encounters a substantial “visual-textual misalignment” problem when applied to TEC-VQA. Specifically, it prioritizes the text in question-answer pairs while disregarding the visual text present in images. Moreover, it fails to address complexities related to nuanced meaning, contextual distortion, language bias, and question-type diversity. In this work, we tackle multilingual TEC-VQA by introducing MTVQA, the first benchmark featuring high-quality human expert annotations across 9 diverse languages, consisting of 6,778 question-answer pairs across 2,116 images. Further, by comprehensively evaluating numerous state-of-the-art Multimodal Large Language Models (MLLMs), including Qwen2-VL, GPT-4o, GPT-4V, Claude3, and Gemini, on the MTVQA benchmark, it is evident that there is still a large room for performance improvement (Qwen2-VL scoring 30.9 versus 79.7 for human performance), underscoring the value of MTVQA. Additionally, we supply multilingual training data within the MTVQA dataset, demonstrating that straightforward fine-tuning with this data can substantially enhance multilingual TEC-VQA performance. We aspire that MTVQA will offer the research community fresh insights and stimulate further exploration in multilingual visual text comprehension.
@@ 1 Introduction
In the era of burgeoning AI, especially in MLLMs [34, 1, 47, 44, 3, 39, 5, 27, 49, 12, 13, 18, 26, 43, 9, 11, 22, 24], Text-Centric Visual Question Answering (TEC-VQA[6, 42, 13, 12, 43, 26, 18] has served as a de facto gold proxy to evaluate AI models in the domain of text-centric scene understanding. Compared with general VQA [6, 30, 36, 42, 32, 31, 29, 52, 20, 4, 28, 40, 25, 15, 14, 23], TEC-VQA places greater emphasis on answering questions that require understanding visual textual information within images. It enables individuals without specialized expertise to access applications in text-centric visual environments. However, most advancements in TEC-VQA have predominantly concentrated on high-resource languages, e.g., English [6, 42, 30, 31], Chinese [37, 15], Japanese [41, 33] and etc., thus restricting the applicability of AI models to the global community, particularly populations speaking low-resource languages.
To tackle the problem of language diversity, several seminal studies [38, 35, 7] in the general VQA field simply leverage translation engines to expand question-answer pairs from high-resource to low-resource languages. However, this translation-based approach is not feasible for TEC-VQA, as it merely processes text in question-answer pairs, neglecting the critical visual text needed for answering based on images. Although the visual text can be recognized by an OCR (Optical Character Recognition) engine and then translated into the target language, this indirect process could lead to a “visual-textual misalignment” problem, due to nuanced meaning, contextual distortion, language bias, and question type diversity. Taking the second case in Fig. 1 as an example, if either the visual text or the question in Russian is recognized or translated problematically, then the question would never be answered correctly. The status quo begs for a question: “Can we directly leverage visual text in source language per se for multilingual TEC-VQA and what we stand in the MLLM era?
@@
In this work, to answer the question above, we establish MTVQA, a novel and high-quality multilingual TEC-VQA benchmark, where all images are collected from real-world and meticulously annotated by human experts in nine languages: Arabic (AR), Korean (KO), Japanese (JA), Thai (TH), Vietnamese (VI), Russian (RU), French (FR), German (DE), and Italian (IT). More concretely, to ensure the visual-textual alignment at most, the annotation process follows the raise-then-correct paradigm, where a group of human annotators raises several distinct questions, ranging from simple content extraction to text-related reasoning, and subsequently provides answers. Another group then double-checks these QA pairs to ensure accuracy and consistency. Consequently, as illustrated in Fig. 2, 6,678 training images and 21,829 question-answer pairs, as well as 2,116 test images and 6,778 question-answer pairs are obtained, covering more than 20 fine-grained scenarios from both documents and natural scenes, such as menus, logos, maps, bills, PPTs, research papers, and etc. To our knowledge, MTVQA is the first TEC-VQA dataset to provide native human annotations for multilingual text-rich scenarios, especially for low-source languages.
Figure 3: Left: comparison of the overall performance of various MLLMs in the MTVQA benchmark. Right: comparison of the performance exhibited by MLLMs in the 9 languages of the MTVQA.
We further investigate recent representative MLLMs in Fig. 3, including GPT-4o, GPT-4V, Gemini, Qwen2-VL, etc., and their performance on MTVQA. As vividly revealed in Fig. 3 (a), the general MLLM Qwen2-VL is the top performer, followed by closed-source GPT-4o and Claude3 Opus. Other general MLLMs like QwenVL Max and QwenVL Plus show mid-tier performance, while text-centric MLLMs lag behind. Fig. 3(b) shows closed-source MLLMs consistently outperforming others, especially in French (FR) and German (DE), while languages like Arabic (AR) and Thai (TH) witness lower scores across all categories. The results unequivocally demonstrate that opportunities for improvement persist within existing MLLMs when applied in multilingual text-rich scenarios. In summary, the main contributions of this paper can be categorized into four points:
@@
  • MTVQA has been serving as a standard benchmark for evaluating multilingual visual text comprehension capabilities by some recently released state-of-the-art MLLMs such as Qwen2-VL [46] and InternVL 2 [11].
@@ 2.1 MLLMs for Text-centric VQA
Recent advancements in MLLMs [1, 47, 44, 3, 39, 5, 27, 49, 12, 13, 18, 26, 43, 9, 11, 22, 24, 51] have revolutionized VQA tasks, as demonstrated by the remarkable zero-shot performance of these models. Notably, the high generalizability of MLLMs, when explicitly trained on visual text understanding datasets and fine-tuned with instructions, has significantly enhanced their application in text-centric VQA scenarios [13, 12, 43, 26, 18]. For example, LLaVAR [50], UniDoc [13], which extend LLaVA [25] into the realm of document understanding, pioneering the text-centric VQA of MLLMs by training them to predict texts and coordinates from document images. Furthermore, DocPedia [12] operates visual input in the frequency domain rather than in space, which enables higher input resolution without increasing the input sequence. Lately, mPLUG-DocOwl [48], Qwen-VL [5], and TextMonkey [26] leverage publicly available document-related VQA datasets to further enhance the text-centric VQA capabilities. Despite the promising results achieved by existing MLLMs in text-centric VQA tasks, their focus on high-resource languages such as English and Chinese has posed challenges in achieving reasonable performance for low-resource languages. This is primarily due to the lack of data or benchmarks for these low-resource languages.
@@ 2.2 Multilingual Text-centric VQA Benchmarks
VQA has garnered significant attention in recent years, with numerous studies, datasets, and benchmarks being proposed to advance the field [6, 30, 36, 42, 32, 31, 29, 52, 20, 4, 28, 40, 25, 15, 14, 23]. Many datasets have been created that encompass scene text of various domains, including natural images [6, 42], scanned documents [30, 31], book and movie covers [32]. One notable limitation of these datasets is their predominant focus on English [6, 42, 30, 31] or other high-resource languages such as Chinese [37, 15] and Japanese [41, 33], which restricts the applicability of VQA systems for low-resource languages such as Thai and Vietnamese.
There are some recent efforts toward extending VQA tasks to a broader range of languages [17, 35, 45, 7, 21, 38] by providing a multilingual VQA datasets. For example,  Gao et al. [15] created a free-form bilingual VQA dataset (FM-IQA) containing over 150,000 images and 310,000 freestyle Chinese question-answer pairs and English translations. Raj Khan et al. [38] developed a large-scale multilingual and code-mixed VQA dataset (MuCo-VQA) supporting five languages. Of more relevance are the works xGQA (7 languages) [35] and MaXM (7 languages) [7], which apply translation-based protocols to expand VQA data beyond English. However, the translation-based multilingual VQA datasets inherently face issues, such as the “visual-textual misalignment” problem, where only the textual information in question-answer pairs is considered, while the visual text in images is overlooked. Additionally, the nuanced meaning and context are often distorted; language bias is introduced by machine translation models, and the coverage of certain question types is limited, as highlighted by [7]. Moreover, none of the previous multilingual datasets focus on text-centric scenarios where multilingual text frequently occurs.
Our benchmark, MTVQA, distinguishes itself by focusing on multilingual text-centric VQA scenarios using human expert annotations. It covers 9 languages, facilitating the training and evaluation of multilingual models in diverse linguistic contexts. Additionally, our dataset can gauge the VQA system’s ability for not only high-resource languages but also those that are typically underrepresented in current datasets [6, 42, 30, 31, 15].
@@ 3.1 Data Collection
The raw data collection aims to gather text-rich images from various scene text and document scenarios, ensuring diversity and quality. This includes images from publicly available datasets (e.g., ICDAR MLT19 [16]) and those sourced from the internet (i.e., Laion-OCR, which is filtered from Common Crawl [10]), such as menus, logos, maps, bills, PowerPoint slides (PPTs), research papers, etc, as shown in Fig. 2. The image collection process consists of two steps: extracting multilingual text using a multilingual OCR engine (Bytedance multilingual OCR API [^1] https://cv-api.bytedance.com/api/ocr-process/v1/multi_language_v2) and selecting by language types and the amount of text contained in the image. Approximately 30% of the overall image data is obtained from public datasets, with 20% sourced from the web and 50% from manual collection, respectively. Manual collection means that the images are taken in different scenarios in the countries and areas where the corresponding language is spoken, which means that the images are native and high quality. A total of 1,220 images from document scenarios and 876 images from natural scenarios are collected for the test set of the MTVQA benchmark. To ensure data content meets regulatory requirements, we subject them to a standardized data cleaning process. The image-cleaning processing pipeline involves a preliminary round of algorithm-driven review to identify and filter out unusable images. It includes detecting and removing images with politically sensitive, pornographic, violent, or other undesirable features. Subsequently, the multilingual OCR tool (Bytedance multilingual OCR API) is employed to extract the textual content from the remaining images. Images devoid of textual information are discarded, and the surviving images are categorized based on their language. Afterward, we organize all the text-rich images we have obtained into language-specific groups, preparing them for the subsequent stage of data annotation.
@@ 3.2 Human Expert Annotation
To obtain informative and accurate text-related QA pairs for images grouped by specific languages, a specialized group of annotators with expertise in the local regions of each language is recruited. The annotators must be native speakers of the corresponding language and have actively utilized it for a minimum duration of 10 years. Additionally, they must possess at least a university-level degree or higher academic qualification, guaranteeing a profound understanding and skillfulness in the linguistic subtleties and cultural contexts required for precise annotations. Given the subjective nature of understanding text within images, the annotation team is divided into two independent groups. One group is tasked with generating questions and providing answers based on the images, while the other group is responsible for evaluating and correcting the QA pairs. This division, known as the raise-then-correct paradigm, ensures a thorough and trustworthy evaluation of the text-rich image comprehension process. Moreover, the annotation results for each language underwent a 10% sampling inspection by a quality inspector to ensure adherence to standards. QA pairs that do not meet the criteria are returned for re-annotation. Prior to the formal annotation process, all annotators will be provided with a detailed explanation of the annotation guidelines, including a dedicated question-and-answer session to clarify any ambiguities and uncertainties. A pilot annotation task will be conducted on a limited dataset to ensure a shared understanding of the annotation rules. The two-round annotation process is briefly illustrated in Fig. 4, further details of which are elaborated in the subsequent sections.
@@
Figure 5: Statistics of question and answer lengths of different languages aggregating training and test set, using GPT-4o tokenizer.
@@ 4.1 Baseline Models
To comprehensively assess MLLMs’ multilingual perception and comprehension capabilities, We select state-of-the-art MLLMs from three categories: open-source general MLLMs, open-source text-centric MLLMs, and closed-source MLLMs. Each category contains the following models: (1) General MLLMs: Qwen2-VL  [46], InternVL-V1.5 [8], InternLM-Xcomposer2-4KHD [11], Mini-Gemini-HD-34B [22], Llava-Next-34B [24], DeepSeek-VL [27], YI-VL-34B [49]; (2) Text-centric MLLMs: TextSquare [43], TextMonkey [26], mPLUG-DocOwl 1.5 [18], MiniCPM-V 2.0 [19]; (3) Closed-source MLLMs: GPT-4o [34], GPT-4V [1], Gemini Ultra [44], QwenVL Max [5], QwenVL Plus [5], Claude3 Opus [3], Claude3 Sonnet [3], and GLM-4V [2]. For the closed-source MLLMs, we use the chat version through official APIs, while for the open-source MLLMs, we utilize the instruct versions available on the HuggingFace Model Hub. The open-source MLLMs’ model size varies from 7b to 76b.
models GPT-4V GPT-4o Gemini Ultra Claude3 Opus Qwen2-VL 92B InternVL2 76B
ANLS 29.6 36.9 31.7 35.4 38.7 29.1
Table 2: Performance of the leading closed- and open-source MLLMs on the MTVQA benchmark. The best results of each language are bolded. The second best results are underlined. “Xcomposer-SFT” denotes instruction tuning to Xcomposer2-4KHD with MTVQA’s training set.
@@
GPT-4V [1] 11.5 31.5 40.4 32.3 11.5 16.7 10.3 15.0 28.9 22.0
GPT-4o [34] 20.2 34.2 41.2 32.7 20.0 33.9 11.5 22.5 34.2 27.8
GPT-4o mini [34] 16.9 33.0 41.2 32.1 18.5 27.4 11.5 19.9 29.1 25.5
Gemini Ultra [44] 14.7 32.3 40.0 31.8 12.3 17.2 11.8 20.3 28.6 23.2
QwenVL Max [5] 7.7 31.4 37.6 30.2 18.6 25.4 10.4 4.8 23.5 21.1
QwenVL Plus [5] 4.8 28.8 33.7 27.1 12.8 19.9 9.4 5.6 18.1 17.8
Claude3 Opus [3] 15.1 33.4 40.6 34.4 19.4 27.2 13.0 19.5 29.1 25.7
Claude3 Sonnet [3] 10.5 28.9 35.6 31.8 13.9 22.2 11.0 15.2 20.8 21.1
GLM-4V [2] 0.3 30.0 34.1 30.1 3.4 5.7 3.0 3.5 12.3 13.6
@@
Qwen2-VL 72B [46] 20.7 36.5 44.1 42.8 21.6 37.4 15.6 17.7 41.6 30.9
InternVL2 76B [1] 9.5 31.3 35.7 35.2 11.1 14.3 11.9 10.0 26.9 22.0
InternVL-V1.5 [8] 3.4 27.1 31.4 27.1 9.9 9.0 4.9 8.7 12.4 14.9
Mini-Gemini-HD-34B [22] 2.2 25.0 29.2 25.5 6.1 8.6 4.1 4.3 11.8 13.0
Llava-Next-34B [24] 3.3 24.0 28.0 22.3 3.6 6.1 2.6 0.4 9.8 11.1
DeepSeek-VL [27] 0.6 14.2 15.3 15.2 2.9 3.8 1.6 0.9 5.2 6.6
YI-VL-34B [49] 1.7 13.5 15.7 12.1 4.8 5.2 0.8 3.5 4.1 6.8
MiniCPM-V 2.0 [19] 1.3 12.7 14.9 17.0 3.7 5.6 2.2 2.2 6.8 7.4
MiniCPM-V 2.5 [19] 6.1 29.6 35.7 26.0 12.1 13.1 5.3 12.6 15.3 17.3
TextSquare [43] 3.7 27.0 30.8 26.7 3.2 7.2 6.7 5.2 12.4 13.6
TextMonkey [26] 2.0 18.1 19.9 22.1 4.6 7.2 3.2 0.9 11.1 9.9
mPLUG-DocOwl 1.5 [18] 1.0 13.9 14.9 18.2 2.9 5.0 2.0 0.9 6.4 7.2
Xcomposer2-4KHD [11] 2.0 20.6 23.2 21.6 5.6 7.7 4.1 6.1 10.1 11.2

### v4

@@ Abstract
Text-Centric Visual Question Answering (TEC-VQA) in its proper format not only facilitates human-machine interaction in text-centric visual environments but also serves as a de facto gold proxy to evaluate AI models in the domain of text-centric scene understanding. Nonetheless, most existing TEC-VQA benchmarks focus on high-resource languages like English and Chinese. Despite pioneering works expanding multilingual QA pairs in non-text-centric VQA datasets through translation engines, the translation-based protocol encounters a substantial “visual-textual misalignment” problem when applied to TEC-VQA. Specifically, it prioritizes the text in question-answer pairs while disregarding the visual text present in images. Moreover, it fails to address complexities related to nuanced meaning, contextual distortion, language bias, and question-type diversity. In this work, we tackle multilingual TEC-VQA by introducing MTVQA, the first benchmark featuring high-quality human expert annotations across 9 diverse languages, consisting of 6,778 question-answer pairs across 2,116 images. Further, by comprehensively evaluating numerous state-of-the-art Multimodal Large Language Models (MLLMs), including Qwen2.5-VL, InternVL-2.5, GPT-4o, GPT-4V, Claude3, and Gemini, on the MTVQA benchmark, it is evident that there is still a large room for performance improvement (InternVL-2.5 scoring 32.2 versus 79.7 for human performance), underscoring the value of MTVQA. By providing a dataset with nuanced multilingual annotations, MTVQA aims to set a new standard for benchmarks, fostering advancements in multilingual visual text comprehension. MTVQA is released at https://github.com/bytedance/MTVQA.
MTVQA: Benchmarking Multilingual Text-Centric Visual Question Answering
Jingqun Tang1, Qi Liu1, Yongjie Ye1, Jinghui Lu1, Shu Wei1, Chunhui Lin1, Wanqing Li1, Mohamad Fitri Faiz Bin Mahmood1, Hao Feng1, Zhen Zhao1, Yanjie Wang1, Yuliang Liu2, Hao Liu1 †, Xiang Bai2, Can Huang1 † 1ByteDance Inc.  2Huazhong University of Science and Technology {tangjingqun, liuqi.nero, yeyongjie.ilz, lujinghui, haoliu.0128, can.huang}@bytedance.com, {ylliu, xbai}@hust.edu.cn
footnotetext: Equal contribution. † Corresponding authors.
@@ 1 Introduction
In the era of burgeoning AI, especially in MLLMs (OpenAI, 2024; Team et al., 2023; Guo et al., 2025; Wang et al., 2025c; Anthropic, 2024; Bai et al., 2023; Lu et al., 2024a; Young et al., 2024; Chen et al., 2024b; Dong et al., 2024; Li et al., 2024; Liu et al., 2024a; Wang et al., 2025a), Text-Centric Visual Question Answering (TEC-VQA(Biten et al., 2019; Singh et al., 2019; Feng et al., 2023b, 2024; Tang et al., 2024; Zhao et al., 2024b, 2025b; Hu et al., 2024; Feng et al., 2025; Lu et al., 2024b, 2025) has served as a de facto gold proxy to evaluate AI models in the domain of text-centric scene understanding. Compared with general VQA, TEC-VQA places greater emphasis on answering questions that require understanding visual text information within images. It enables individuals without specialized expertise to access applications in text-centric visual environments. However, most advancements in TEC-VQA have predominantly concentrated on high-resource languages, e.g., English (Biten et al., 2019; Singh et al., 2019; Mathew et al., 2021, 2022; Wang et al., 2025b), Chinese (Qi et al., 2022; Gao et al., 2015; Fu et al., 2024), Japanese (Shimizu et al., 2018; Nguyen et al., 2023), and etc., thus restricting the applicability of AI models to the global community, particularly populations speaking low-resource languages.
To tackle the problem of language diversity, several seminal studies (Raj Khan et al., 2021; Pfeiffer et al., 2022; Changpinyo et al., 2023) in the general VQA field simply leverage translation engines to expand question-answer pairs from high-resource to low-resource languages. However, this translation-based approach is not feasible for TEC-VQA, as it merely processes text in question-answer pairs, neglecting the critical visual text needed for answering based on images. Although the visual text can be recognized by an OCR (Optical Character Recognition) (Zhao et al., 2024a; Tang et al., 2022c, b, a, 2023; Liu et al., 2023) engine and then translated into the target language, this indirect process could lead to a “visual-textual misalignment” problem, due to nuanced meaning, contextual distortion, language bias, and question type diversity. Taking the second case in Fig. 1 as an example, if either the visual text or the question in Russian is recognized or translated problematically, then the question would never be answered correctly. The status quo begs for a question: “Can we directly leverage visual text in source language per se for multilingual TEC-VQA and what we stand in the MLLM era?
@@
In this work, to answer the question above, we establish MTVQA, a novel and high-quality multilingual TEC-VQA benchmark, where all images are collected from real-world and meticulously annotated by human experts in nine languages: Arabic (AR), Korean (KO), Japanese (JA), Thai (TH), Vietnamese (VI), Russian (RU), French (FR), German (DE), and Italian (IT). More concretely, to ensure the visual-textual alignment at most, the annotation process follows the maker-checker paradigm, where a group of human annotators raises several distinct questions, ranging from simple content extraction to text-related reasoning, and subsequently provides answers. Another group then double-checks these QA pairs to ensure accuracy and consistency. Consequently, as illustrated in Fig. 2, 6,678 training images and 21,829 question-answer pairs, as well as 2,116 test images and 6,778 question-answer pairs are obtained, covering more than 20 fine-grained scenarios from both documents and natural scenes, such as menus, logos, maps, bills, PPTs, research papers, and etc. To our knowledge, MTVQA is the first TEC-VQA dataset to provide native human annotations for multilingual text-rich scenarios, especially for low-resource languages.
We further investigate recent representative MLLMs, including GPT-4o, GPT-4V, Gemini, Qwen2.5-VL, InternVL-2.5, etc., and their performance on MTVQA. As vividly revealed in Tab. 2, the general MLLM InternVL-2.5 is the top performer, followed by Qwen2.5-VL and closed-source GPT-4o and Claude3 Opus. Other general MLLMs like QwenVL Max and QwenVL Plus show mid-tier performance, while text-centric MLLMs lag behind. The results unequivocally demonstrate that opportunities for improvement persist within existing MLLMs when applied in multilingual text-rich scenarios. In summary, the main contributions of this paper can be categorized into four points:
@@ 2.1 MLLMs for Text-centric VQA
Recent advancements in MLLMs have revolutionized VQA tasks, as demonstrated by the remarkable zero-shot performance of these models. Notably, the high generalizability of MLLMs, when explicitly trained on visual text understanding datasets and fine-tuned with instructions  (Lu et al., 2025; Fei et al., 2025; Ke et al., 2025; Qiu et al., 2025; Zhao et al., 2025a; Ouyang et al., 2024; Shen and Tang, 2024; Shen et al., 2025), has significantly enhanced their application in text-centric VQA scenarios (Feng et al., 2023b, a; Tang et al., 2024; Liu et al., 2024c; Hu et al., 2024). For example, LLaVAR (Zhang et al., 2023), UniDoc (Feng et al., 2023b), which extend LLaVA (Liu et al., 2024b) into the realm of document understanding, pioneering the text-centric VQA of MLLMs by training them to predict texts and coordinates from document images. Furthermore, DocPedia (Feng et al., 2023a) operates on visual input in the frequency domain rather than in space, which enables higher input resolution without increasing the input sequence. Lately, mPLUG-DocOwl (Ye et al., 2023), Qwen-VL (Bai et al., 2023), and TextMonkey (Liu et al., 2024c) leverage publicly available document-related VQA datasets to further enhance the text-centric VQA capabilities. What’s more, TextHarmony (Zhao et al., 2024b) unifies visual text generation and comprehension in a single model. Despite the promising results achieved by existing MLLMs in text-centric VQA tasks (He et al., 2025; Sun et al., 2025; Shan et al., 2024; Wang et al., 2025c), their focus on high-resource languages such as English and Chinese has posed challenges in achieving reasonable performance for low-resource languages. This is primarily due to the lack of data or benchmarks for these low-resource languages.
@@ 2.2 Multilingual Text-centric VQA Benchmarks
VQA has garnered significant attention in recent years, with numerous studies, datasets, and benchmarks being proposed to advance the field. Many datasets have been created that encompass scene text of various domains, including natural images (Biten et al., 2019; Singh et al., 2019), scanned documents (Mathew et al., 2021, 2022), books, and movie covers (Mishra et al., 2019). One notable limitation of these datasets is their predominant focus on English (Biten et al., 2019; Singh et al., 2019; Mathew et al., 2021, 2022) or other high-resource languages such as Chinese (Qi et al., 2022; Gao et al., 2015) and Japanese (Shimizu et al., 2018; Nguyen et al., 2023), which restricts the applicability of VQA systems for low-resource languages such as Thai and Vietnamese.
There are some recent efforts toward extending VQA tasks to a broader range of languages (Gupta et al., 2020; Pfeiffer et al., 2022; Vivoli et al., 2022; Changpinyo et al., 2023; Li et al., 2023; Raj Khan et al., 2021) by providing multilingual VQA datasets. For example,  Gao et al. (2015) created a free-form bilingual VQA dataset (FM-IQA) containing over 150,000 images and 310,000 freestyle Chinese question-answer pairs and English translations. Raj Khan et al. (2021) developed a large-scale multilingual and code-mixed VQA dataset (MuCo-VQA) supporting five languages. Of more relevance are the works xGQA (7 languages) (Pfeiffer et al., 2022) and MaXM (7 languages) (Changpinyo et al., 2023), which apply translation-based protocols to expand VQA data beyond English. However, the translation-based multilingual VQA datasets inherently face issues, such as the “visual-textual misalignment” problem, where only the textual information in question-answer pairs is considered, while the visual text in images is overlooked. Additionally, the nuanced meaning and context are often distorted; language bias is introduced by machine translation models, and the coverage of certain question types is limited, as highlighted by Changpinyo et al. (2023). Moreover, none of the previous multilingual datasets focus on text-centric scenarios where multilingual text frequently occurs.
@@ 3.1 Data Collection
The raw data collection aims to gather text-rich images from various scene text and document scenarios, ensuring diversity and quality. This includes images from publicly available datasets (e.g., ICDAR MLT19 (Gao et al., 2019)) and those sourced from the internet (i.e., Laion-OCR, which is filtered from Common Crawl (Crawl, 2024)), such as menus, logos, maps, bills, PowerPoint slides (PPTs), research papers, etc, as shown in Fig. 2. The image collection process consists of two steps: extracting multilingual text using a multilingual OCR engine (Bytedance multilingual OCR APIhttps://cv-api.bytedance.com/api/ocr-process/v1/multi_language_v2) and selecting by language types and the amount of text contained in the image. Approximately 30% of the overall image data is obtained from public datasets, with 20% sourced from the web and 50% from manual collection, respectively. Manual collection means that the images are taken in different scenarios in the countries and areas where the corresponding language is spoken, which means that the images are native and high quality. A total of 1,220 images from document scenarios and 876 images from natural scenarios are collected for the test set of the MTVQA benchmark. To ensure data content meets regulatory requirements, we subject it to a standardized data cleaning process. The image-cleaning processing pipeline involves a preliminary round of algorithm-driven review to identify and filter out unusable images. It includes detecting and removing images with politically sensitive, pornographic, violent, or other undesirable features. Subsequently, the multilingual OCR tool (Bytedance multilingual OCR API) is employed to extract the textual content from the remaining images. Images devoid of textual information are discarded, and the surviving images are categorized based on their language. Afterward, we organize all the text-rich images we have obtained into language-specific groups, preparing them for the subsequent stage of data annotation.
@@ 3.2 Human Expert Annotation
To obtain informative and accurate text-related QA pairs for images grouped by specific languages, a specialized group of annotators with expertise in the local regions of each language is recruited. The annotators must be native speakers of the corresponding language and have actively utilized it for a minimum duration of 10 years. Additionally, they must possess at least a university-level degree or higher academic qualification, guaranteeing a profound understanding and skillfulness in the linguistic subtleties and cultural contexts required for precise annotations. Given the subjective nature of understanding text within images, the annotation team is divided into two independent groups. One group is tasked with generating questions and providing answers based on the images, while the other group is responsible for evaluating and correcting the QA pairs. This division, known as the maker-checker paradigm, ensures a thorough and trustworthy evaluation of the text-rich image comprehension process. Moreover, the annotation results for each language underwent a 10% sampling inspection by a quality inspector to ensure adherence to standards. QA pairs that do not meet the criteria are returned for re-annotation. Prior to the formal annotation process, all annotators will be provided with a detailed explanation of the annotation guidelines, including a dedicated question-and-answer session to clarify any ambiguities and uncertainties. A pilot annotation task will be conducted on a limited dataset to ensure a shared understanding of the annotation rules. The two-round annotation process is briefly illustrated in Fig. 3, further details of which are elaborated in the subsequent sections.
@@
Figure 4: Statistics of question and answer lengths of different languages aggregating training and test sets, using GPT-4o tokenizer.
@@ 4.1 Baseline Models
To comprehensively assess MLLMs’ multilingual perception and comprehension capabilities, we select state-of-the-art MLLMs from three categories: open-source general MLLMs, open-source text-centric MLLMs, and closed-source MLLMs.
@@
GPT-4V (Achiam et al., 2023) 11.5 31.5 40.4 32.3 11.5 16.7 10.3 15.0 28.9 22.0
GPT-4o (OpenAI, 2024) 20.2 34.2 41.2 32.7 20.0 33.9 11.5 22.5 34.2 27.8
GPT-4o mini (OpenAI, 2024) 16.9 33.0 41.2 32.1 18.5 27.4 11.5 19.9 29.1 25.5
Gemini Ultra (Team et al., 2023) 14.7 32.3 40.0 31.8 12.3 17.2 11.8 20.3 28.6 23.2
QwenVL Max (Bai et al., 2023) 7.7 31.4 37.6 30.2 18.6 25.4 10.4 4.8 23.5 21.1
QwenVL Plus (Bai et al., 2023) 4.8 28.8 33.7 27.1 12.8 19.9 9.4 5.6 18.1 17.8
Claude3 Opus (Anthropic, 2024) 15.1 33.4 40.6 34.4 19.4 27.2 13.0 19.5 29.1 25.7
Claude3 Sonnet (Anthropic, 2024) 10.5 28.9 35.6 31.8 13.9 22.2 11.0 15.2 20.8 21.1
GLM-4V (AI, 2024) 0.3 30.0 34.1 30.1 3.4 5.7 3.0 3.5 12.3 13.6
@@
Qwen2-VL 72B (Wang et al., 2024) 20.7 36.5 44.1 42.8 21.6 37.4 15.6 17.7 41.6 30.9
Qwen2.5-VL 72B (Qwen, 2025) 17.2 34.2 38.5 35.2 18.4 29.6 13.9 9.5 41.6 28.7
InternVL2 76B (Chen et al., 2023) 9.5 31.3 35.7 35.2 11.1 14.3 11.9 10.0 26.9 22.0
InternVL2.5 78B  (Chen et al., 2024a) 15.9 39.0 45.6 42.9 21.1 33.9 12.2 23.8 41.5 32.2
Mini-Gemini-HD-34B (Li et al., 2024) 2.2 25.0 29.2 25.5 6.1 8.6 4.1 4.3 11.8 13.0
Llava-Next-34B (Liu et al., 2024a) 3.3 24.0 28.0 22.3 3.6 6.1 2.6 0.4 9.8 11.1
DeepSeek-VL (Lu et al., 2024a) 0.6 14.2 15.3 15.2 2.9 3.8 1.6 0.9 5.2 6.6
YI-VL-34B (Young et al., 2024) 1.7 13.5 15.7 12.1 4.8 5.2 0.8 3.5 4.1 6.8
MiniCPM-V 2.0 (Hu et al., 2023) 1.3 12.7 14.9 17.0 3.7 5.6 2.2 2.2 6.8 7.4
MiniCPM-V 2.5 (Hu et al., 2023) 6.1 29.6 35.7 26.0 12.1 13.1 5.3 12.6 15.3 17.3
TextSquare (Tang et al., 2024) 3.7 27.0 30.8 26.7 3.2 7.2 6.7 5.2 12.4 13.6
TextMonkey (Liu et al., 2024c) 2.0 18.1 19.9 22.1 4.6 7.2 3.2 0.9 11.1 9.9
mPLUG-DocOwl 1.5 (Hu et al., 2024) 1.0 13.9 14.9 18.2 2.9 5.0 2.0 0.9 6.4 7.2
Xcomposer2-4KHD (Dong et al., 2024) 2.0 20.6 23.2 21.6 5.6 7.7 4.1 6.1 10.1 11.2
@@
Table 2: Performance of the leading closed- and open-source MLLMs on the MTVQA benchmark. The best results of each language are bolded. The second-best results are underlined. “Xcomposer-SFT” denotes instruction tuning to Xcomposer2-4KHD with MTVQA’s training set.